# Can a calibration metric be both testable and actionable?

### v1

@@ D.2 Modifying Platt scaling to achieve Cutoff calibration
Define the usual Platt-scaling post-processor as $\tilde{h}_{n}$. Then, our final $h_{n}$ is designed to ensure $\widehat{\Delta}_{\operatorname{Cutoff}}(h_{n}\circ f)$ is small, up to a user-specified threshold $\epsilon_{n}$.
$h_{n}(z)=\begin{cases}\tilde{h}_{n}(z)&\text{if }\widehat{\Delta}_{\operatorname{Cutoff}}(\tilde{h}_{n}\circ f)\leq\epsilon_{n}\\ \mathbb{E}_{n}[Y]&\text{otherwise}\end{cases}$ (114)
@@
Suppose $(X_{i},Y_{i})\overset{iid}{\sim}P$ for $i=1,\dots,n$. Assume $Y_{i}\in[0,1]$ and $f$ is fixed. Let $h_{n}$ be as defined in (114). Fix $\delta>0$. Let $\epsilon_{n}>0$ be a user-specified tolerance for model-misspecification. Then, with probability at least $1-\delta$,
$\Delta_{\operatorname{Cutoff}}^{P}(h_{n}\circ f)\leq\frac{30+2\sqrt{2\log(2/\delta)}}{\sqrt{n}}+\epsilon_{n}$ (115)
@@
$\widehat{\Delta}_{\operatorname{Cutoff}}(\mathbb{E}_{n}[Y])=0$ (116)
@@
$\widehat{\Delta}_{\operatorname{Cutoff}}(h_{n}\circ f)\leq\epsilon_{n}$ (117)
@@
$\Delta_{\operatorname{Cutoff}}^{P}(h_{n}\circ f)$ (118)
$=\Delta_{\operatorname{Cutoff}}^{P}(h_{n}\circ f)-\widehat{\Delta}_{\operatorname{Cutoff}}(h_{n}\circ f)+\widehat{\Delta}_{\operatorname{Cutoff}}(h_{n}\circ f)$ (119)
$\leq\Delta_{\operatorname{Cutoff}}^{P}(h_{n}\circ f)-\widehat{\Delta}_{\operatorname{Cutoff}}(h_{n}\circ f)+\epsilon_{n}$ (120)
$\leq\sup_{I:\text{ interval}}\left|\mathbb{E}[(Y-h_{n}\circ f(X)){{\mathbf{1}_{\left\{h_{n}\circ f(X)\in I\right\}}}}]-\frac{1}{n}\sum_{i=1}^{n}(Y_{i}-h_{n}\circ f(X_{i})){{\mathbf{1}_{\left\{h_{n}\circ f(X_{i})\in I\right\}}}}\right|+\epsilon_{n}$ (121)
The first term can be shown to be $\lesssim\frac{1}{\sqrt{n}}$ with probability at least $1-\delta$ using the same analysis as we did for isotonic regression, since $h_{n}$ is also monotone here. ∎

### v2

@@ D.2 Modifying Platt scaling to achieve Cutoff calibration
Define the usual Platt-scaling post-processor as $\tilde{h}_{n}$. Then, our final $h_{n}$ is designed to ensure $\widehat{\Delta}_{\operatorname{Cutoff}}(h_{n}\circ f)$ is small, up to a user-specified threshold $\epsilon_{n}$:
$h_{n}(z)=\begin{cases}\tilde{h}_{n}(z)&\text{if }\widehat{\Delta}_{\operatorname{Cutoff}}(\tilde{h}_{n}\circ f)\leq\epsilon_{n}\\ \mathbb{E}_{n}[Y]&\text{otherwise}\end{cases}.$
@@
Suppose $(X_{i},Y_{i})\overset{iid}{\sim}P$ for $i=1,\dots,n$. Assume $Y_{i}\in[0,1]$ and $f$ is fixed. Let $h_{n}$ be as defined in (D.2). Fix $\delta>0$. Let $\epsilon_{n}>0$ be a user-specified tolerance for model-misspecification. Then, with probability at least $1-\delta$,
$\Delta_{\operatorname{Cutoff}}^{P}(h_{n}\circ f)\leq\frac{30+2\sqrt{2\log(2/\delta)}}{\sqrt{n}}+\epsilon_{n}.$
@@
    $\widehat{\Delta}_{\operatorname{Cutoff}}(\mathbb{E}_{n}[Y])=0.$
@@
    $\widehat{\Delta}_{\operatorname{Cutoff}}(h_{n}\circ f)\leq\epsilon_{n}.$
@@
    $\Delta_{\operatorname{Cutoff}}^{P}(h_{n}\circ f)$
    $=\Delta_{\operatorname{Cutoff}}^{P}(h_{n}\circ f)-\widehat{\Delta}_{\operatorname{Cutoff}}(h_{n}\circ f)+\widehat{\Delta}_{\operatorname{Cutoff}}(h_{n}\circ f)$
    $\leq\Delta_{\operatorname{Cutoff}}^{P}(h_{n}\circ f)-\widehat{\Delta}_{\operatorname{Cutoff}}(h_{n}\circ f)+\epsilon_{n}$
    $\leq\sup_{I:\text{ interval}}\left|\mathbb{E}[(Y-h_{n}\circ f(X)){{\mathbf{1}_{\left\{h_{n}\circ f(X)\in I\right\}}}}]-\frac{1}{n}\sum_{i=1}^{n}(Y_{i}-h_{n}\circ f(X_{i})){{\mathbf{1}_{\left\{h_{n}\circ f(X_{i})\in I\right\}}}}\right|+\epsilon_{n}.$
    The first term can be shown to be $\lesssim\frac{1}{\sqrt{n}}$ with probability at least $1-\delta$ following the same arguments in the isotonic regression proof, from equation (5) onward, since $h_{n}$ is also monotone here. ∎